Nov 25, 2020 Calorimeter
Meeting Day: WEDNESDAY
Meeting Time: 11:00 a.m. JLab time
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone in CC L207 is 757-269-7084 (usual room)
- JLab Phone: in CC F326 is 757-269-6460
- Phone in the Regina Video-conference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
- FCAL HDFCAL log book
- BCAL HDBCAL log book
- Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
Goals for Calorimetry Group
- Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
- Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
- Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.
- Run Preparations
- BCAL Gain Calibration for Summer 2020 Data (B9-B12).
- BCAL LED
- BCAL (offline)
- FCAL (offline) bad channels map GlueX data
- Quick task force simulation update/question (Tolga) 
- Z Calibration Study update Update
- Action Items
- Any other business
Attending: Karthik, Varun, Tolga, Rebecca, Mark, Sean, Sasha, Elton, Susan, Zisis
- Announcements: it's Thanksgiving tomorrow!
- Run Preparations
- FCAL: Mark reported that repairs to the first 100 bases (center portion of the detector) have been completed. Testing is pending. On the firmware side, the radiation test did not work. Ben wants to send a signal to the base to lock it up. More bases will be fixed in the new year.
- BCAL: no activity.
- Karthik reported on the BCAL Gain Calibration for Summer 2020 Data (B9-B12). Two iterations have been complete on B9 and things look nominal. Elton asked how long will it take to complete the calibration. Karthik said that it should be done by Monday.
- BCAL LED: no activity
- BCAL (offline): no offline monitoring activity at present.
- FCAL (offline) Susan presented an update on the bad-channels map. Main issue is HV stability and the result will provide channel efficiencies for each run. The LEDs fall into 5 regions (0-4); regions can be used to lock onto which LED-bank is firing. The Blue 15V LED was used for this work. LED skims were analyzed using a plugin adapted from PrimEx/Chandra. All physics production runs of RunPeriod 2018-08 were analyzed, and the results are available on the ifarm. Responding to a question by Sasha, Mark said that the table block_mc_efficiency in the ccdb is used (block_quality table is not used). The idea is to process all the runs from 2017 through 2018. For 2017 data, this will be more involved because there are no LED regions and the gains were being adjusted. Zisis suggested DocDB-2195 (Athens) as a reference. Zisis/Sean mentioned that evio files can be merged or multiple ones can be passed on the command line.
- Tolga showed slides with his starting tests on the tag and probe method (gp->p eta, eta-g(g). Plots included photon momentum versus angle; gaps appear (e.g. near 24 deg) that need to be understood. Elton suggested to look at the missing (g). The discussion will continue offline to check whether both gammas are shown in these plots. Elton will reconvene the reconstruction efficiency task force to discuss this topic.
- Karthik picked up the Z CalibrationsStudy update, since it was last looked at in the spring. He is using bggen and sees an issue with E_raw and the Z distribution, both of which have spikes at -1 and 220 when using DANAREST. This is not the case when using smeared files. But smeared files were not saved for the entire/large bggen data set that Rebecca generated. A pi0 gun shows the same behaviour. A discussion took place on whether to update the skim plugin and save skims one bggen data, or skim evio files vs hddm ones. Sean will help Karthik investigate.
- Simulations: nothing to report.
- Action Items: no discussion
- Any other business: none